Please solve.
(t-1)^2=7

Answer by stanbon(48564)

(t-1)^2=7
Take the square root of both sides to get:
t-1 = sqrt(7) or t-1 = -sqrt(7)
t = 1+sqrt(7) or t = 1-sqrt(7)
=============
Cheers,
Stan H.